OpenStudy (anonymous):

Drew is making a cardboard cut out of a right triangle. He knows that the longest side must be 15 inches long. One of the short sides must be 8 inches. How many inches long will he make the other side? Use the appropriate formula to find the solution and round to the nearest tenth. Show all steps or reasoning used. Be sure to check your answer for reasonableness

OpenStudy (lgbasallote):

right triangle so use pythagorean formula...longest side means hypotenuse so \[a^2 + (8^2) = 15^2\]

OpenStudy (lgbasallote):

therefore \[a^2 = 15^2 - 8^2\] to get a just take the square root \[a = \sqrt{15^2 - 8^2}\] got ti??

OpenStudy (anonymous):

sqrt161

OpenStudy (lgbasallote):

do you get what i did @Isuckatmath1995 ??

OpenStudy (anonymous):

yess

Parth (parthkohli):

Good work, what do you get as the answer?

OpenStudy (lgbasallote):

so i assume you can do the rest?

OpenStudy (lgbasallote):

i was talking to @Isuckatmath1995 btw @ParthKohli not you

Parth (parthkohli):

As I said, \( \color{Black}{\Rightarrow \sqrt{15^2 - 8^2} }\) First find out what \(15^2\) and \(8^2\) are.

OpenStudy (lgbasallote):

i gtg @Isuckatmath1995 so just input that in the calculator to find your answer ^_^ bye

OpenStudy (anonymous):

thnaks for your welcome :)

OpenStudy (anonymous):

*your help lol

OpenStudy (lgbasallote):

you're welcome :) haha

OpenStudy (anonymous):

:D